# Rajamanickam, Almutari and Majumdar

### v2

@@ 6 Conclusions
We have rigorously explored the influence of colloidal spherical NPs on cholesteric liquid crystals using the homogenized LdG framework as developed in [7], in the dilute limit. For spatially homogeneous systems, the LdG bulk energy $f_{B}$ in (5) is replaced by the modified bulk potential, $f_{m}$ in (10), to account for the collective effects of suspended spherical NPs with identical boundary treatments, in the dilute limit. The suspended NPs are fully characterised by $W$ (the anchoring strength of the NP–LC interface), the NP geometry (spherical in this case), and the preferred boundary condition, $\mathbf{Q}_{\nu}$, on all NP-LC interfaces. The preferred boundary condition is imposed by means of a Rapini–Papoular surface energy [7], although other choices are possible too. The modified bulk potential $f_{m}$ differs from $f_{B}$ through an effective temperature $\tilde{\pazocal{A}}$ and a linear energy contribution induced by a constant matrix $\mathbf{X}$, that originates from the homogenized contribution of the NP-LC surface energies summed over all NP-LC interfaces. The effective temperature increases for $W>0$, and the bulk symmetry is broken if $\mathbf{X}\neq 0$. Notably, we consider different possibilities for $\mathbf{X}$, corresponding to arbitrary $\mathbf{Q}_{\nu}$; in this respect, our study is perhaps the most general to date. If $\mathbf{X}=0$, then the critical points of $f_{m}$ are either uniaxial or isotropic (as in the case of $f_{B}$) but the critical temperatures are shifted, i.e., we get a first-order isotropic-nematic phase transition at a lower temperature for $W>0$, as compared with the $W=0$ case. For $\mathbf{X}\neq 0$, we consider four representative cases: (i) uniaxial $\mathbf{X}$ with prolate or oblate symmetry and (ii) biaxial $\mathbf{X}$ with prolate or oblate symmetry. In all cases, we obtain seven critical points of $f_{m}$ which approach the critical points of $f_{B}$ for low temperatures (i.e., as $\pazocal{A}\to$-$\infty$). The global minimizer of $f_{m}$ always has prolate symmetry at low temperatures, as in the case of $f_{B}$. Furthermore, we always observe locally stable biaxial bulk phases induced by $\mathbf{X}$, for a certain range of temperatures, which are otherwise inaccessible without the NPs. There is no first order isotropic-nematic phase transition when $\mathbf{X}\neq 0$ and we instead observe a gradual crossover between ordered and disordered states, with increasing temperature. At high temperatures, there is a unique minimizer of $f_{m}$ that inherits the prolate or oblate symmetry of $\mathbf{X}$, again a behaviour that is not possible without the NPs. Our work thus highlights the capacity of NPs to stabilize phases that are otherwise inaccessible in pure nematic or cholesteric systems.
@@
We complement our work on spatially homogeneous samples with a simple example of a dilute suspension of spherical NPs in a cholesteric-filled channel geometry, subject to planar Dirichlet boundary conditions on the the top and bottom surfaces. This is a well-studied example and people expect to find stable twisted cholesteric director profiles, when the channel is sufficiently large, i.e., when the confinement parameter $\lambda$ is sufficiently large. However, this symmetry is broken by the suspended NPs, with increasing temperature. We demonstrate structural transitions from stable cholesteric helices to almost uniform director profiles, with increasing temperature, provided $\lambda$ is sufficiently large. The uniform director profiles inherit the symmetry of $\mathbf{X}$ when $\pazocal{A}$ and $\lambda$ are large enough. Equally importantly, this transition is reversible, i.e., we can numerically observe untwisted - twisted helical transitions induced by slowly decreasing the temperature, for sufficiently large $\lambda$. While this is a specific example, the qualitative conclusions will remain unchanged for generic confined cholesteric or nematic systems. Such systems tend to exhibit multistablility in large geometries (i.e., with large $\lambda$) under suitably prescribed boundary conditions. In other words, there are typically multiple LdG energy minimizers corresponding to multiple physically observable functional modes of the large LC system. This multistability will typically be lost when $\mathbf{X}\neq 0$, since the NPs impose a unique director orientation (specified by the symmetry of $\mathbf{X}$), for sufficiently large $\pazocal{A}$ and $\lambda$. An especially interesting example would be a cholesteric or nematic-filled three-dimensional spherical droplet with radial boundary conditions [20, 28], which naturally hosts interior defects, as a consequence of non-trivial topology. The radial symmetry can be broken by the suspended NPs in a spherical droplet, if $\mathbf{X}\neq 0$, and it would be interesting to investigate how $\mathbf{X}$ tailors interior defect structures in such confined systems in future work.
@@
Overall, our findings underscore the potential of NPs as a versatile tool to engineer novel phases and structural transitions in cholesteric liquid crystals. We believe that our conclusions will remain valid for general NP shapes (other than spheres) and also for generic surface energies, within the dilute limit. For example, in [7, 8], the authors consider a generic surface energy density on the NP-LC interface given by
@@ Acknowledgments
A.M. gratefully acknowledges support from the Leverhulme Research Project Grant RPG-2021-401 and an Isaac Newton Institute Network Grant, EPSRC Grant EP/R014604/1. P.R. is supported by the Leverhulme Research Project Grant RPG-2021-401.

### v3

@@ 6 Conclusions
We have rigorously explored the influence of spherical colloidal NPs on liquid crystals using the homogenized LdG framework as developed in [7], in the dilute limit. For spatially homogeneous systems, the LdG bulk energy $f_{B}$ in (5) is replaced by the modified bulk potential, $f_{m}$ in (10), to account for the collective effects of suspended spherical NPs with identical boundary treatments, in the dilute limit. The suspended NPs are fully characterised by $W$ (the anchoring strength of the NP–LC interface), the NP geometry (spherical in this case), and the preferred boundary condition, $\mathbf{Q}_{\nu}$, on all NP-LC interfaces. The preferred boundary condition is imposed by means of a Rapini–Papoular surface energy [7], although other choices are possible too. The modified bulk potential $f_{m}$ differs from $f_{B}$ through an effective temperature $\tilde{\pazocal{A}}$ and a linear energy contribution induced by a constant matrix $\mathbf{X}$, that originates from the homogenized contribution of the NP-LC surface energies summed over all NP-LC interfaces. The effective temperature increases for $W>0$, and the bulk symmetry is broken if $\mathbf{X}\neq 0$. Notably, we consider different possibilities for $\mathbf{X}$, corresponding to arbitrary $\mathbf{Q}_{\nu}$; in this respect, our study is perhaps the most general to date. If $\mathbf{X}=0$, then the critical points of $f_{m}$ are either uniaxial or isotropic (as in the case of $f_{B}$) but the critical temperatures are shifted, i.e., we get a first-order isotropic-nematic phase transition at a lower temperature for $W>0$, as compared with the $W=0$ case. For $\mathbf{X}\neq 0$, we consider four representative cases: (i) uniaxial $\mathbf{X}$ with prolate or oblate symmetry and (ii) biaxial $\mathbf{X}$ with prolate or oblate symmetry. In all cases, we obtain seven critical points of $f_{m}$ which approach the critical points of $f_{B}$ for low temperatures (i.e., as $\pazocal{A}\to$-$\infty$). The global minimizer of $f_{m}$ always has prolate symmetry at low temperatures, as in the case of $f_{B}$. Furthermore, we always observe locally stable biaxial bulk phases induced by $\mathbf{X}$, for a certain range of temperatures, which are otherwise inaccessible without the NPs. There is no first order isotropic-nematic phase transition when $\mathbf{X}\neq 0$ and we instead observe a gradual crossover between ordered and disordered states, with increasing temperature. At high temperatures, there is a unique minimizer of $f_{m}$ that inherits the prolate or oblate symmetry of $\mathbf{X}$, again a behaviour that is not possible without the NPs. Our work thus highlights the capacity of NPs to stabilize phases that are otherwise inaccessible in pure nematic or cholesteric systems.
@@
We complement our work on spatially homogeneous samples with a simple example of a dilute suspension of spherical NPs in a cholesteric-filled channel geometry, subject to planar Dirichlet boundary conditions on the the top and bottom surfaces. This is a well-studied example and people expect to find stable twisted cholesteric director profiles, when the channel is sufficiently large, i.e., when the confinement parameter $\lambda$ is sufficiently large. However, this symmetry is broken by the suspended NPs, with increasing temperature. We demonstrate structural transitions from stable cholesteric helices to almost uniform director profiles with increasing temperature, provided $\lambda$ is sufficiently large. The uniform director profiles inherit the symmetry of $\mathbf{X}$ when $\pazocal{A}$ and $\lambda$ are large enough. Equally importantly, this transition is reversible, i.e., we can numerically observe untwisted - twisted helical transitions induced by slowly decreasing the temperature, for sufficiently large $\lambda$. While this is a specific example, the qualitative conclusions will remain unchanged for generic confined cholesteric or nematic systems. Such systems tend to exhibit multistablility in large geometries (i.e., with large $\lambda$) under suitably prescribed boundary conditions. In other words, there are typically multiple LdG energy minimizers corresponding to multiple physically observable functional modes of the large LC system. This multistability will typically be lost when $\mathbf{X}\neq 0$, since the NPs impose a unique director orientation (specified by the symmetry of $\mathbf{X}$), for sufficiently large $\pazocal{A}$ and $\lambda$. An especially interesting example would be a cholesteric or nematic-filled three-dimensional spherical droplet with radial boundary conditions [19, 27], which naturally hosts interior defects, as a consequence of non-trivial topology. The radial symmetry can be broken by the suspended NPs in a spherical droplet, if $\mathbf{X}\neq 0$, and it would be interesting to investigate how $\mathbf{X}$ tailors interior defect structures in such confined systems in future work.
@@
Overall, our findings underscore the potential of NPs as a versatile tool to engineer novel phases and structural transitions in liquid crystals. We believe that our conclusions will remain valid for general NP shapes (other than spheres) and also for generic surface energies, within the dilute limit. For example, in [7, 8], the authors consider a generic surface energy density on the NP-LC interface given by
@@ Acknowledgments
A.M. gratefully acknowledges support from the Leverhulme Research Project Grant RPG-2021-401 and an Isaac Newton Institute Network Grant, EPSRC Grant EP/R014604/1. P.R. is supported by the Leverhulme Research Project Grant RPG-2021-401. This work was completed whilst A.M. was working at the University of Strathclyde.